# Non-Hermitian Hamiltonian Approach for Two-Dimensional Spectroscopy

### v1

@@ 4 Conclusion
In this paper, we compare the NHH method and the RF formalism with their advantages and disadvantages in the obtaining of 2DS. In particular, we give a simple simplification to the computation processes of the NHH method by introducing the quasi-Green function, which improves its interpretive capacity. Both methods are demonstrated by gaining the RP signal of 2DS for the three-level system under the influence of a control field. The NHH signal is decomposed into the sum of the products, of which three quasi-Green functions corresponding to different $t_{c}$ are multiplied. An intuitive explanation of the NHH method is thus given directly in terms of the Feynman diagrams. It is worth noting that the NHH method yields signal with more Liouville paths and evolution information. We compare the population dynamics of the two methods in $t_{2}$, obtaining similar oscillatory behavior in the first few periods, but different steady states for the long-time evolution. The comparison of the sum of populations on the three levels in $t_{2}$ indicates that for the RF formalism the probability is conserved, while for the NHH method the probability keeps decaying. This decay leads to the different steady states in the population dynamics, and also to a weakening of the emitted signal as $t_{2}$ increases. To sum up, although the NHH method suffers from the limitation of non-conservative population, it is particularly suitable for investigating dissipative or complex systems. Besides, the equation-of-motion method, which has a conserved population but an even higher computational complexity, may help in this scene 51.
@@
We plot the 2DS for four representative $t_{2}$’s. Both methods yield splitting peaks with the same locations for each split peak. For the RF formalism, the heights of the four peaks oscillate synchronously. In contrast, for the NHH method, the four peaks oscillate for a few periods with continuous attenuation, and the oscillation amplitude of the non-diagonal peaks are more significant than that of the diagonal peaks. We show that the disappearance and emergence of off-diagonal peaks in 2DS by the NHH approach are attributed to the interference of different Feynman diagrams induced by the control fields. Furthermore, by comparing the 2DS of the two methods with the same Liouville paths, we show that the main difference between the two approaches is due to the introduction of some additional Feynman diagrams and the disappearance of some less-important Feynman diagrams induced by the control fields.
@@
Overall, the differences between the 2DS of the two methods are attributed to the fact that in the NHH method there red are unidirectional population losses to the environment, and there are more abundant Liouville paths. The NHH method has its advantage of obtaining the signals directly from the time evolution, and thus is less likely to omit any possible Liouville paths. This approach can be more applicable in complicated systems with complex energy structure and transition paths, as compared to the RF formalism. It will help to obtain 2DS more comprehensively and has prospects for exploring the dynamics of molecular or atomic systems with complex protocols.
@@
We thank Lipeng Chen for valuable comments on the paper. This work is supported by Innovation Program for Quantum Science and Technology under grant No. 2023ZD0300200, the Natural Science Foundation of Beijing Municipality under Grant No. 1202017 and the National Natural Science Foundation of China under Grant Nos. 11674033, 11505007, and Beijing Normal University under Grant No. 2022129.

### v2

@@ 4 Conclusion
In this paper, we compare the NHH method and the RF formalism with their advantages and disadvantages in the obtaining of 2DCS. In particular, we propose a simple way to simplify the computation processes of the NHH method by introducing the quasi-Green function, which improves its interpretive capacity. Both methods are demonstrated by gaining the RP signal of 2DCS for the same three-level system under the influence of a control field. The NHH signal is decomposed into the sum of the products, of which three quasi-Green functions corresponding to different $t_{c}~{}(c=1,2,3)$ are multiplied. An intuitive explanation of the NHH method is thus given directly in terms of the Feynman diagrams. It is worth noting that the NHH method yields signal with more Liouville paths and evolution information. We compare the population dynamics of the two methods in $t_{2}$, obtaining similar oscillatory behavior in the first few periods, but different steady states for the long-time evolution. The comparison of the sum of populations on the three levels in $t_{2}$ indicates that for the RF formalism the probability is conserved, while for the NHH method the probability keeps decaying. This decay leads to the different steady states in the population dynamics, and also to a weakening of the emitted signal as $t_{2}$ increases. On this issue, the equation-of-motion method is capable of handling these scenarios despite its higher computational complexity 79, i.e., $O(N^{2})$ vs $O(N)$, with $N$ being the number of states. This method is powerful since it does not require the use of Feynman diagrams. Moreover, it can be combined with thermo field dynamics to account for more realistic temperature effects. 80, 81, 82
@@
We plot the 2DCS for four representative $t_{2}$’s. Both methods yield splitting peaks with the same locations for each split peak. For the RF formalism, the heights of the four peaks oscillate synchronously. In contrast, the four peaks of the NHH method oscillate for a few periods with continuous attenuation, and the oscillation amplitudes of the non-diagonal peaks vary more significantly than those of the diagonal peaks. We show that the disappearance and emergence of off-diagonal peaks in 2DCS by the NHH approach are attributed to the different Feynman diagrams induced by the control field. Furthermore, by comparing the 2DCS of the two methods with the same NHH Liouville paths, we show that the main difference between the two approaches is due to the introduction of some additional Feynman diagrams and the disappearance of some less-important Feynman diagrams induced by the control fields. The RF formalism in its original form is simply inapplicable to driven systems with time-dependent Hamiltonians. Though it can be applied if we switch to the rotating frame with a time-independent Hamiltonian. In addition to the simplified model system, we simulate the 2DCS of RDC dissolved in hexane. The positions and signs of all peaks are faithfully reproduced by the NHH approach, and is in good agreement with the experimental plot. Therefore, by comparing the 2DCS from experiments with those predicted by the NHH approach, we can effectively extract useful information about the structure and dynamics of the molecule under investigation.
Compared to the Hamiltonian in the RF formalism, the NHH includes the relaxation but drops the quantum-jump terms. Meanwhile, the RF formalism takes the relaxation into account by solving the quantum master equation without dropping these terms. This approximation in the NHH method leads to a non-conserved trace in population dynamics, and manifests as an overall decay in the intensity of the 2DCS. However, the overall decay does not significantly influence other phenomena, such as the oscillations in the off-diagonal peaks. Despite the overall decay, other differences arise from the method adopted, with the most significant being the disappearance and reappearance of off-diagonal peaks.
@@
Overall, the main differences between the 2DCS of the two methods arise from the fact that, in the NHH method, the relaxations are equivalent to unidirectional population losses to the environment. Additionally, there are more significant Liouville paths contributing to the final 2DCS. The NHH method has its advantage of obtaining the signals directly from the analytical solution of time evolution, and thus would not omit any dominant Liouville paths. However, since we assume constant decay rates, the NHH method is not suited for lineshape analysis and thus can not account for homogeneous broadening. As for the RF formalism, its results are reliable as long as it considers all the important Liouville paths. To sum up, the NHH method can be more applicable in the systems with abundant energy structure and transition paths, as compared to the RF formalism. It will help to obtain 2DCS with all dominant Liouville paths considered and has prospects for exploring the dynamics of molecular or atomic systems with control fields.
@@
We thank Lipeng Chen for valuable comments on the paper. This work is supported by Innovation Program for Quantum Science and Technology under Grant No. 2023ZD0300200, the National Natural Science Foundation of China under Grant No. 62461160263, Beijing Natural Science Foundation under Grant No. 1202017, and Beijing Normal University under Grant No. 2022129.